A boy throwns a stone horizontally off a cliff. it hits the sea 2 seconds later,at a distance of 40 m from the foot of the cliff.    Calculate the                                                                             (a) height of the cliff                                                                   (b) intial speed of the stone                                                      © direction in which the stone is moving when it strikes the water.

Grade:12th pass

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Height = 0.5 g t*t
= 0.5*10*2*2= 20

Initial speed = horizontal dist/ time

= 40/ 2 = 20 m/s scalar quantity

Direction of stone is in the direction of initial speed if no obstacle is present

Regards